SELF- AND AIR-BROADENED OXYGEN A-BAND PARAMETERS

\underline{L. R. BROWN} AND C. E. MILLER,, Jet Propulsion Laboratory,California Institute of Technology, Pasadena, CA 91109; A. S. PINE, Technologies, 14401 Poplar Hill Road, Germantown, MD 20874.

Self- and air-broadened oxygen A-band spectra, previously reported by Brown and Plymate^{a} were combined with new higher pressure (1 to 3 atm) spectra of air and fitted with a multispectrum analysis to investigate lineshapes and to improve the precision of the molecular parameters for the O_{2} A-band at 13100 cm^{−1}. These data were recorded with the McMath-Pierce Fourier-transform interferometer located at the National Solar Observatory on Kitt Peak in Arizona. The transition wavenumbers have been referenced to recent stabilized cavity-ring-down laser measurements of Robichaud et al.^{b} calibrated against a ^{39}K standard. Deviations from the conventional Voigt profile were observed while better fits were obtained with Dicke-narrowing (Rautian or Galatry) and speed-dependent Voigt linehapes. Line-mixing, apparent in the higher pressure air spectra, was successfully fit through the R-branch head with first-order Rosenkranz coefficients exhibiting odd-even J staggering indicative of interbranch spin-rotation coupling. The results will be compared to other literature measurements and theoretical models of the oxygen A-band.^{c}

aL. R. Brown and C. Plymate, J. Mol. Spectrosc. 199, 166 (2000).

bD. J. Robichaud, J. T. Hodges, P. Maslowski, L. Y. Yeung, M. Okumura, C. E Miller and L. R. Brown, J. Mol. Spectrosc.

(2008) in press

cPart of the research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.